# Proton Decay and Gauge Coupling Unification in an Extended SU(5) GUT with 45-Dimensional Higgs

### v1

@@ 2 SU(5) GUT with 45 dimensional Higgs
$\bm{\bar{5}}:\overline{\psi}^{a}_{i}=\begin{pmatrix}d_{Ri}^{Ca}&\epsilon_{\alpha\beta}\,(V_{DL})_{i}{}^{j}\,\ell_{Lj}^{\,\beta}\end{pmatrix},~{}~{}~{}~{}\bm{10}:\chi_{i\,ab}=\frac{1}{\sqrt{2}}\begin{pmatrix}\epsilon^{abc}\,(V_{QU})_{i}{}^{j}\,u^{C}_{Rj\,c}&q_{Li}^{\,\hat{a}\beta}\\[2.84526pt] -q_{Li}^{\,b\alpha}&\epsilon^{\alpha\beta}\,(V_{QE})_{i}{}^{j}\,e_{Rj}^{\,c}\end{pmatrix}.$ (1)
Here, $a,b,c$ are SU(3) indices, $\alpha,\beta$ are SU(2) indices, and $i,j$ are generation indices. The unitary matrices $V_{DL}$, $V_{QU}$, and $V_{QE}$ are determined by imposing matching conditions between the GUT and SM fermion sectors at the unification scale.
@@
$-\mathcal{L}_{Y}$ $=\frac{1}{8}\,(Y_{5}^{U})_{ij}\epsilon^{abcde}\,\chi_{i\,ab}\,\chi_{j\,cd}\,(5_{H})_{e}+(Y_{5}^{D})_{ij}\chi_{i\,ab}\,\overline{\psi}^{a}_{i}\,(5_{H}^{\dagger})^{b}$
$+\frac{1}{4}(Y_{45}^{U})_{ij}\epsilon^{abcde}\,\chi_{i\,ab}\,\chi_{i\,cf}\,({45}_{H})_{de}^{f}+\frac{1}{2}\,(Y_{45}^{D})_{ij}\chi_{i\,ab}\,\overline{\psi}^{c}_{j}({45}_{H}^{\dagger})_{ab}^{c}+\mathrm{h.c.}.$ (2)
@@
where $c_{H}=\cos\theta_{H}$, $s_{H}=\sin\theta_{H}$, $c_{S}=\cos\theta_{S}$, and $s_{S}=\sin\theta_{S}$. The SU(5) invariant Yukawa interactions in Eq.(2) can be decomposed in terms of the component fields as follows [14]:
@@
$+\frac{(Y_{C}^{QQ})_{ij}}{2}\,\epsilon_{abc}\,\epsilon_{\alpha\beta}\,\bar{q}^{C\,a\,\alpha}_{Li}H_{C}^{*\,bq_{Lj}^{c}\,\beta}+\frac{(Y_{1}^{QQ})_{ij}}{2}\,\epsilon_{\hat{a}\hat{b}\hat{c}}\,\epsilon_{\alpha\beta}\,\bar{q}^{\,c\,\hat{a}\alpha}_{Li}S_{1}^{*\hat{b}}q_{Lj}^{\hat{c}\beta}$
$+(Y_{1}^{QL})_{ij}\epsilon_{\alpha\beta}\,\bar{q}^{\,c\,\hat{a}\alpha}_{Li}S_{1\hat{a}}\,\ell_{Lj}^{\,\beta}+(Y_{1}^{UE})_{ij}\,\bar{u}_{R\hat{a}i}\,S_{1}^{*\hat{a}}e_{Rj}^{c}+(Y_{1}^{DU})_{ij}\epsilon^{\hat{a}\hat{b}\hat{c}}\,\bar{d}_{R\hat{a}i}\,S_{1\hat{b}}\,u_{R\hat{c}j}^{c}$
$+(\tilde{Y}_{1}^{ED})_{ij}\,\bar{e}_{Ri}\,\tilde{S}_{1}^{*\hat{a}}d_{R\hat{a}j}^{\,c}+\frac{(\tilde{Y}_{1}^{UU})_{ij}}{2}\,\epsilon^{\hat{a}\hat{b}\hat{c}}\,\bar{u}_{R\hat{a}i}\,\tilde{S}_{1\hat{b}}\,u_{R\hat{c}j}^{c}+(Y_{2}^{UL})_{ij}\epsilon_{\alpha\beta}\,\bar{u}_{R\hat{a}i}\,R_{2}^{\,\hat{a}\alpha}\ell_{Lj}^{\,\beta}+(Y_{2}^{EQ})_{ij}\,\bar{e}_{Ri}\,R_{2\hat{a}\alpha}^{\,*}\,q_{Lj}^{\hat{a}\alpha}$
$+(Y_{3}^{QL})_{ij}\epsilon_{\alpha\beta}\,\bar{q}_{Li}^{\,c\,\hat{a}\gamma}(\sigma_{a})^{\alpha}{}_{\gamma}\,S_{3\hat{a}}^{a}\,\ell_{Lj}^{\,\beta}+\frac{(Y_{3}^{QQ})_{ij}}{2}\,\epsilon_{\hat{a}\hat{b}\hat{c}}\,\epsilon_{\alpha\beta}\,\bar{q}_{Li}^{c\,\hat{a}\alpha}(\sigma^{a})^{\beta}{}_{\gamma}\,S_{3}^{*a\hat{b}}q_{Lj}^{\hat{c}\,\gamma}$
$+(Y_{6}^{DU})_{ij}\,\bar{d}_{R\hat{a}i}\,(\eta^{A})^{\hat{a}\hat{b}}S_{6}^{A}\,u_{R\hat{b}j}^{c}+\frac{(Y_{6}^{QQ})_{ij}}{2}\,\epsilon_{\alpha\beta}\,\bar{q}_{Li}^{\,c\,\hat{a}\alpha}(\eta^{A})_{\hat{a}\hat{b}}\,S_{6}^{A*}\,q_{Lj}^{\hat{b}\beta}$
$+(Y_{8}^{UQ})_{ij}\epsilon_{\alpha\beta}\,\bar{u}_{R\hat{a}i}\,(\lambda^{A})^{\hat{a}}{}_{\hat{b}}\,S_{8}^{A\alpha}\,q_{Lj}^{\hat{b}\beta}+(Y_{8}^{DQ})_{ij}\,\bar{d}_{R\hat{a}i}\,(\lambda^{A})^{\hat{a}}{}_{\hat{b}}\,S_{8\alpha}^{A*}\,q_{Lj}^{\hat{b}\alpha}+\mathrm{h.c.}\,,$ (6)
@@
where $\hat{Y}_{U}$, $\hat{Y}_{D}$, and $\hat{Y}_{E}$ represent diagonal matrices in the mass basis and $V_{\text{CKM}}$ is the Cabbibo-Kobayashi-Maskawa matrix, which are determined by the observed fermion masses and mixing angles. On the other hand, the unitary matrices, $V_{QE}$, $V_{QU}$, $V_{DL}$, cannot be determined solely from low-energy experiments. These matrices represent the rotations required to diagonalize the quark-Higgs and lepton-Higgs Yukawa couplings in the SU(5) GUT framework and are important for the proton decay analysis.
@@ 3 Gauge Coupling Unification
$\alpha_{2}^{-1}(M_{Z})=\alpha_{2}^{-1}(\Lambda)-\frac{1}{2\pi}\biggl{[}$ $-\frac{19}{6}\log\frac{M_{Z}}{\Lambda}-\frac{21}{2}\log\frac{M_{XY}}{\Lambda}+\frac{1}{3}\log\frac{M_{\Sigma}}{\Lambda}+\frac{1}{6}\log\frac{M_{S_{1}^{(45)}}}{\Lambda}$
@@
$\alpha_{2}^{-1}(M_{Z})-3\alpha_{2}^{-1}(M_{Z})+2\alpha_{3}^{-1}(M_{Z})=-\frac{1}{2\pi}\biggl{[}$ $\frac{2}{5}\log\frac{M_{H_{C}}}{M_{Z}}-\frac{2}{5}\log\frac{M_{H^{(45)}}}{\Lambda}+\frac{2}{5}\log\frac{M(\overline{3},1)}{\Lambda}$
@@
$-5\alpha_{2}^{-1}(M_{Z})+3\alpha_{2}^{-1}(M_{Z})+2\alpha_{3}^{-1}(M_{Z})=-\frac{1}{2\pi}\biggl{[}$ $-42\log\frac{M_{XY}}{M_{Z}}-2\log\frac{M_{\Sigma}}{M_{Z}}+6\log\frac{M_{S_{3}}}{\Lambda}$
@@ 4 Proton decay
To suppress the proton decay rate mediated by the $S_{3}$ component and satisfy the experimental constraints, we adopt the assumption that the GUT Yukawa coupling $Y^{U}_{45}$ is forbidden at the GUT scale, following the approach in Ref. [14]. Even with this assumption, it is possible to reproduce realistic fermion masses and mixings within the framework of the extended GUT model. We focus our analysis on the proton decay process mediated by the exchange of the $S_{1}$ component. The partial decay widths for these processes are calculated using the effective Lagrangian and taking into account the relevant hadronic matrix elements and form factors as follows [15]:
@@
where $Y_{1}^{DU}$, $Y^{QL}$, $Y^{UE}_{1}$ are the Yukawa coupling matrices in the mass basis. These mixing parameters are evaluated at the energy scale $M_{S_{1}}$ by solving the corresponding RGEs.
@@
To numerically calculate the proton decay rates, we follow a systematic procedure. First, we solve the RGEs for the SM Yukawa couplings, evolving them from the electroweak scale up to the GUT scale, $\Lambda=6.5\times 10^{15}~{}\mathrm{GeV}$. This GUT scale is determined by the reference values of the scalar masses given in Eq. (16), which lead to successful GCU. For the input parameters at the electroweak scale, we use the values listed in Table 1 [23]. By solving the RGEs, we obtain the values of the Yukawa coupling matrices $Y_{U}$, $Y_{D}$, and $Y_{E}$ at the GUT scale. From these Yukawa coupling matrices at the GUT scale, we can calculate the GUT-scale Yukawa couplings $Y^{U}_{5}$, $Y^{D}_{5}$, $Y^{U}_{45}$, and $Y^{D}_{45}$ using the matching equations in Eq. (8). Importantly, we assume that the Yukawa coupling $Y^{U}_{45}$ vanishes at the GUT scale, following the assumption made in the previous discussion. This assumption implies that the unitary matrix $V_{QE}$ should be the identity matrix to satisfy the matching conditions. For the other unitary matrices, $V_{QU}$ and $V_{DL}$, we assume that they are random unitary matrices whose elements are distributed according to the Haar measure as in Ref. [24], because their specific forms are not directly constrained by low-energy observables. We also assume that $\delta_{H}$ and $\theta_{H}$ are random values. Once the GUT-scale Yukawa couplings are determined, we can obtain the Yukawa couplings for the light scalar components of the 45-Higgs, namely $S_{8}$, $S_{3}$, and $S_{1}$, from the matching conditions in Eqs. (7). With these boundary conditions at the GUT scale, we solve the RGEs for the Yukawa couplings, including the contributions from the light scalar components of $45_{H}$, evolving them from the GUT scale down to the mass scale of $S_{1}$. Using the Yukawa couplings at the $M_{S_{1}}$ scale, we can calculate the partial decay widths for various proton decay modes, such as $p\rightarrow\mu^{+}K^{0}$ and $p\rightarrow\mu^{+}\pi^{0}$, using the formulae given in Eqs. (19)-(24). These partial decay widths depend on the Yukawa couplings, the scalar masses, and other input parameters related to the hadronic matrix elements and form factors. If we further evolve these Yukawa couplings from the $M_{S_{1}}$ scale down to the electroweak scale using the RGEs, the resulting Yukawa coupling matrices $Y_{U}$, $Y_{D}$, and $Y_{E}$ would differ from their initial input values at the electroweak scale. This discrepancy arises because we first evolved the Yukawa couplings using the SM RGEs from the electroweak scale to the GUT scale, and then introduced additional contributions from the scalar components of $45_{H}$ when evolving from the GUT scale to the $M_{S_{1}}$ scale. While this discrepancy exists, it is not expected to significantly affect the estimation of the proton decay rates, which primarily depend on the Yukawa couplings at the $M_{S_{1}}$ scale. For simplicity and to avoid unnecessary complications, we neglect this discrepancy and focus on the Yukawa couplings at the $M_{S_{1}}$ scale when calculating the proton decay rates.
@@
In Fig. 2, we present correlations among the lower bounds on the mass $M_{S_{1}}$ derived from the experimental constraints on various proton decay modes. We observe that the data points in the plot between $p\to\nu\pi$ and $p\to\mu K$ modes exhibit some scatter. This scatter can be attributed to the different contributions to the decay widths of these two modes. The decay width for the $p\to\nu\pi$ mode is primarily determined by the term $|\delta^{L}|^{2}$, which represents the contribution from the left-handed mixing parameters $\delta^{L}_{ijkl}$. On the other hand, the decay width for the $p\to\mu K$ mode receives contributions from both $|\delta^{L}|^{2}$ and $|\delta^{R}|^{2}$ terms. The $|\delta^{R}|^{2}$ term represents the contribution from the right-handed mixing parameters $\delta^{R}_{ijkl}$. The presence of both $|\delta^{L}|^{2}$ and $|\delta^{R}|^{2}$ contributions in the $p\to\mu K$ decay width leads to a larger spread in the values of the lower bound on $M_{S_{1}}$ compared to the $p\to\nu\pi$ mode, which depends only on $|\delta^{L}|^{2}$. This difference in the dependence on the mixing parameters results in the observed scatter between the data points in the plot. These plots demonstrate the importance of properly constraining these parameters using experimental data from different proton decay channels to obtain robust bounds on the mass scale $M_{S_{1}}$.

### v2

@@ 2 SU(5) GUT with 45 dimensional Higgs
$\bm{\bar{5}}:\overline{\psi}_{\hat{a}i}=\begin{pmatrix}d_{Rai}^{C}&\epsilon_{\alpha\beta}\,(V_{DL})_{i}{}^{j}\,\ell_{Lj}^{\,\beta}\end{pmatrix},~{}~{}~{}~{}\bm{10}:\chi_{i}^{\hat{a}\hat{b}}=\frac{1}{\sqrt{2}}\begin{pmatrix}\epsilon^{abc}\,(V_{QU})_{i}{}^{j}\,u^{C}_{Rcj}&q_{Li}^{\,a\beta}\\[2.84526pt] -q_{Li}^{\,b\alpha}&\epsilon^{\alpha\beta}\,(V_{QE})_{i}{}^{j}\,e_{Rj}^{\,C}\end{pmatrix}.$ (1)
Here, $\hat{a},\hat{b}$ are SU(5) indices, $a,b,c$ are SU(3) indices, $\alpha,\beta$ are SU(2) indices, and $i,j$ are generation indices. The unitary matrices $V_{DL}$, $V_{QU}$, and $V_{QE}$ are determined by imposing matching conditions between the GUT and SM fermion sectors at the unification scale.
@@
$-\mathcal{L}_{Y}$ $=\frac{1}{8}\,(Y_{5}^{U})_{ij}\epsilon_{\hat{a}\hat{b}\hat{c}\hat{d}\hat{e}}\,\chi_{i}^{\hat{a}\hat{b}}\,\chi_{j}^{\hat{c}\hat{d}}\,(5_{H})^{\hat{e}}+(Y_{5}^{D})_{ij}\chi_{i}^{\hat{a}\hat{b}}\,\overline{\psi}_{\hat{a}j}\,(5_{H}^{\dagger})_{\hat{b}}$
$+\frac{1}{4}(Y_{45}^{U})_{ij}\epsilon_{\hat{a}\hat{b}\hat{c}\hat{d}\hat{e}}\,\chi_{i}^{\hat{a}\hat{b}}\,\chi_{i}^{\hat{c}\hat{f}}\,({45}_{H})^{\hat{d}\hat{e}}_{\hat{f}}+\frac{1}{2}\,(Y_{45}^{D})_{ij}\chi_{i}^{\hat{a}\hat{b}}\,\overline{\psi}_{\hat{c}j}({45}_{H}^{\dagger})_{\hat{a}\hat{b}}^{\hat{c}}+\mathrm{h.c.}.$ (2)
@@
where $c_{H}=\cos\theta_{H}$, $s_{H}=\sin\theta_{H}$, $c_{S}=\cos\theta_{S}$, and $s_{S}=\sin\theta_{S}$. [^1] We have assumed that the colored Higgs in the 5-dimensional Higgs boson, $S_{1}^{(5)}(3,1)_{-\frac{1}{3}}$, has a mass around the GUT scale and the lighter mass eigenstate, $S_{1}$, is primarily composed of $S_{1}^{(45)}(3,1)_{-\frac{1}{3}}$ with a tiny mixing angle $\theta_{S}\sim m^{2}_{S_{1}^{(45)}}/m^{2}_{S_{1}^{(5)}}\ll 1$. The SU(5) invariant Yukawa interactions in Eq.(2) can be decomposed in terms of the component fields as follows [14]:
@@
$+\frac{(Y_{C}^{QQ})_{ij}}{2}\,\epsilon_{abc}\,\epsilon_{\alpha\beta}\,\bar{q}^{C\,a\,\alpha}_{Li}H_{C}^{*\,b}q_{Lj}^{C}\,\beta+\frac{(Y_{1}^{QQ})_{ij}}{2}\,\epsilon_{abc}\,\epsilon_{\alpha\beta}\,\bar{q}^{\,C\,a\alpha}_{Li}S_{1}^{*\hat{b}}q_{Lj}^{C\beta}$
$+(Y_{1}^{QL})_{ij}\epsilon_{\alpha\beta}\,\bar{q}^{\,C\,a\alpha}_{Li}S_{1a}\,\ell_{Lj}^{\,\beta}+(Y_{1}^{UE})_{ij}\,\bar{u}_{Rai}\,S_{1}^{*a}e_{Rj}^{C}+(Y_{1}^{DU})_{ij}\epsilon^{abc}\,\bar{d}_{Rai}\,S_{1b}\,u_{Rcj}^{C}$
$+(\tilde{Y}_{1}^{ED})_{ij}\,\bar{e}_{Ri}\,\tilde{S}_{1}^{*a}d_{Raj}^{\,C}+\frac{(\tilde{Y}_{1}^{UU})_{ij}}{2}\,\epsilon^{abc}\,\bar{u}_{Rai}\,\tilde{S}_{1b}\,u_{Rcj}^{C}+(Y_{2}^{UL})_{ij}\epsilon_{\alpha\beta}\,\bar{u}_{Rai}\,R_{2}^{\,a\alpha}\ell_{Lj}^{\,\beta}+(Y_{2}^{EQ})_{ij}\,\bar{e}_{Ri}\,R_{2a\alpha}^{\,*}\,q_{Lj}^{a\alpha}$
$+(Y_{3}^{QL})_{ij}\epsilon_{\alpha\beta}\,\bar{q}_{Li}^{\,C\,a\gamma}(\sigma_{A})^{\alpha}{}_{\gamma}\,S_{3a}^{A}\,\ell_{Lj}^{\,\beta}+\frac{(Y_{3}^{QQ})_{ij}}{2}\,\epsilon_{abc}\,\epsilon_{\alpha\beta}\,\bar{q}_{Li}^{C\,a\alpha}(\sigma_{A})^{\beta}{}_{\gamma}\,S_{3}^{*Ab}q_{Lj}^{C\,\gamma}$
$+(Y_{6}^{DU})_{ij}\,\bar{d}_{Rai}\,(\eta^{A})^{ab}S_{6}^{A}\,u_{Rbj}^{C}+\frac{(Y_{6}^{QQ})_{ij}}{2}\,\epsilon_{\alpha\beta}\,\bar{q}_{Li}^{\,C\,a\alpha}(\eta^{A})_{ab}\,S_{6}^{A*}\,q_{Lj}^{b\beta}$
$+(Y_{8}^{UQ})_{ij}\epsilon_{\alpha\beta}\,\bar{u}_{Rai}\,(\lambda^{A})^{a}{}_{b}\,S_{8}^{A\alpha}\,q_{Lj}^{b\beta}+(Y_{8}^{DQ})_{ij}\,\bar{d}_{Rai}\,(\lambda^{A})^{a}{}_{b}\,S_{8\alpha}^{A*}\,q_{Lj}^{b\alpha}+\mathrm{h.c.}\,,$ (6)
@@
where $\hat{Y}_{U}$, $\hat{Y}_{D}$, and $\hat{Y}_{E}$ represent diagonal matrices in the mass basis and $V_{\text{CKM}}$ is the Cabibbo-Kobayashi-Maskawa matrix, which are determined by the observed fermion masses and mixing angles. On the other hand, the unitary matrices, $V_{QE}$, $V_{QU}$, $V_{DL}$, cannot be determined solely from low-energy experiments. These matrices represent the rotations required to diagonalize the quark-Higgs and lepton-Higgs Yukawa couplings in the SU(5) GUT framework and are important for the proton decay analysis.
@@ 3 Gauge Coupling Unification
$\alpha_{2}^{-1}(M_{Z})=\alpha_{2}^{-1}(\Lambda)-\frac{1}{2\pi}\biggl{[}$ $-\frac{19}{6}\log\frac{M_{Z}}{\Lambda}-\frac{21}{2}\log\frac{M_{XY}}{\Lambda}+\frac{1}{3}\log\frac{M_{\Sigma}}{\Lambda}+\frac{1}{6}\log\frac{M_{S_{1}}}{\Lambda}$
@@
$\alpha_{1}^{-1}(M_{Z})-3\alpha_{2}^{-1}(M_{Z})+2\alpha_{3}^{-1}(M_{Z})=-\frac{1}{2\pi}\biggl{[}$ $\frac{2}{5}\log\frac{M_{H_{C}}}{M_{Z}}-\frac{2}{5}\log\frac{M_{H^{(45)}}}{\Lambda}+\frac{2}{5}\log\frac{M_{S_{1}}}{\Lambda}$
@@
$-5\alpha_{1}^{-1}(M_{Z})+3\alpha_{2}^{-1}(M_{Z})+2\alpha_{3}^{-1}(M_{Z})=-\frac{1}{2\pi}\biggl{[}$ $-42\log\frac{M_{XY}}{M_{Z}}-2\log\frac{M_{\Sigma}}{M_{Z}}+6\log\frac{M_{S_{3}}}{\Lambda}$
@@ 4 Proton decay
To suppress the proton decay rate mediated by the $S_{3}$ component and satisfy the experimental constraints, we adopt the assumption that the GUT Yukawa coupling $Y^{U}_{45}$ is forbidden at the GUT scale, following the approach in Ref. [14]. Even if $Y^{U}_{45}$ is forbidden at the GUT scale, $Y_{3}^{QQ}$ can be generated through a one-loop RGE effect involving $S_{1}$. However, the corresponding term of the one-loop beta function of $Y_{3}^{QQ}$ is proportional to the product $(Y_{1}^{QL})(Y_{3}^{QL})^{*}(Y_{1}^{QQ})$. Since we have assumed that $S_{1}$ is predominantly composed of the 45-dimensional Higgs boson, $Y_{1}^{QQ}$ is highly suppressed by the tiny mixing angle $\theta_{S}$. Consequently, the one-loop contribution to $Y_{3}^{QQ}$ is negligible, ensuring that the proton decay rate remains sufficiently suppressed. Even with this assumption, it is possible to reproduce realistic fermion masses and mixings within the framework of the extended GUT model. We focus our analysis on the proton decay process mediated by the exchange of the $S_{1}$ component. The partial decay widths for these processes are calculated using the effective Lagrangian and taking into account the relevant hadronic matrix elements and form factors as follows [15]:
@@
where $Y_{1}^{DU}$, $Y_{1}^{QL}$, $Y^{UE}_{1}$ are the Yukawa coupling matrices in the mass basis. These mixing parameters are evaluated at the energy scale $M_{S_{1}}$ by solving the corresponding RGEs. Notice that the proton decay mediated by the $Y_{1}^{QQ}$ coupling is negligible since the $S_{1}$ scalar is predominantly composed of the 45-dimensional Higgs boson and $Y_{1}^{QQ}$ is highly suppressed by the tiny mixing angle $\theta_{S}$.
@@
To numerically calculate the proton decay rates, we follow a systematic procedure. First, we solve the RGEs for the SM Yukawa couplings, evolving them from the electroweak scale up to the GUT scale, $\Lambda=6.5\times 10^{15}~{}\mathrm{GeV}$. This GUT scale is determined by the reference values of the scalar masses given in Eq. (16), which lead to successful GCU. For the input parameters at the electroweak scale, we use the values listed in Table 1 [16]. By solving the RGEs, we obtain the values of the Yukawa coupling matrices $Y_{U}$, $Y_{D}$, and $Y_{E}$ at the GUT scale. From these Yukawa coupling matrices at the GUT scale, we can calculate the GUT-scale Yukawa couplings $Y^{U}_{5}$, $Y^{D}_{5}$, $Y^{U}_{45}$, and $Y^{D}_{45}$ using the matching equations in Eq. (8). Importantly, we assume that the Yukawa coupling $Y^{U}_{45}$ vanishes at the GUT scale, following the assumption made in the previous discussion. This assumption implies that the unitary matrix $V_{QU}$ should be the identity matrix to satisfy the matching conditions. For the other unitary matrices, $V_{QE}$ and $V_{DL}$, we assume that they are random unitary matrices whose elements are distributed according to the Haar measure as in Ref. [18], because their specific forms are not directly constrained by low-energy observables. We also assume that $\delta_{H}$ and $\theta_{H}$ are random values. Once the GUT-scale Yukawa couplings are determined, we can obtain the Yukawa couplings for the light scalar components of the 45-Higgs, namely $S_{8}$, $S_{3}$, and $S_{1}$, from the matching conditions in Eqs. (7). With these boundary conditions at the GUT scale, we solve the RGEs for the Yukawa couplings, including the contributions from the light scalar components of $45_{H}$, evolving them from the GUT scale down to the mass scale of $S_{1}$. Using the Yukawa couplings at the $M_{S_{1}}$ scale, we can calculate the partial decay widths for various proton decay modes, such as $p\rightarrow\mu^{+}K^{0}$ and $p\rightarrow\mu^{+}\pi^{0}$, using the formulae given in Eqs. (19)-(24). These partial decay widths depend on the Yukawa couplings, the scalar masses, and other input parameters related to the hadronic matrix elements and form factors. If we further evolve these Yukawa couplings from the $M_{S_{1}}$ scale down to the electroweak scale using the RGEs, the resulting Yukawa coupling matrices $Y_{U}$, $Y_{D}$, and $Y_{E}$ would differ from their initial input values at the electroweak scale. This discrepancy arises because we first evolved the Yukawa couplings using the SM RGEs from the electroweak scale to the GUT scale, and then introduced additional contributions from the scalar components of $45_{H}$ when evolving from the GUT scale to the $M_{S_{1}}$ scale. While this discrepancy exists, it is not expected to significantly affect the estimation of the proton decay rates, which primarily depend on the Yukawa couplings at the $M_{S_{1}}$ scale. For simplicity and to avoid unnecessary complications, we neglect this discrepancy and focus on the Yukawa couplings at the $M_{S_{1}}$ scale when calculating the proton decay rates.
@@
In Fig. 2, we present correlations among the lower bounds on the mass $M_{S_{1}}$ derived from the experimental constraints on various proton decay modes. We observe that the data points in the plot between $p\to\nu\pi$ and $p\to\mu K$ modes exhibit some scatter. This scatter can be attributed to the different contributions to the decay widths of these two modes. The decay width for the $p\to\nu\pi$ mode is primarily determined by the term $|\delta^{L}|^{2}$, which represents the contribution from the left-handed mixing parameters $\delta^{L}_{ijkl}$. On the other hand, the decay width for the $p\to\mu K$ mode receives contributions from both $|\delta^{L}|^{2}$ and $|\delta^{R}|^{2}$ terms. The $|\delta^{R}|^{2}$ term represents the contribution from the right-handed mixing parameters $\delta^{R}_{ijkl}$. The presence of both $|\delta^{L}|^{2}$ and $|\delta^{R}|^{2}$ contributions in the $p\to\mu K$ decay width leads to a larger spread in the values of the lower bound on $M_{S_{1}}$ compared to the $p\to\nu\pi$ mode, which depends only on $|\delta^{L}|^{2}$. Furthermore, the left-handed mixing parameter for $p\to\mu K$, $\delta^{L}_{2112}=(Y_{1}^{DU})^{*}_{21}(Y^{QL}_{1})_{12}$, involves the product of off-diagonal elements from two Yukawa matrices. This contrasts with other decay modes, where at least one element is diagonal. As off-diagonal elements tend to have a wider spread than diagonal elements, the $p\to\mu K$ mode exhibits the most significant scatter among all decay modes. This difference in the dependence on the mixing parameters results in the observed scatter between the data points in the plot. These plots demonstrate the importance of properly constraining these parameters using experimental data from different proton decay channels to obtain robust bounds on the mass scale $M_{S_{1}}$.